Jack is 3 years younger than Martin. Rob is 7 years older than Jack. The total age of the three boys is 25. Find the age of each boy.
Answer by macston(5194)

J=Jack's age; M=Marin's age=J+3; R=Rob's age=J+7
.
J+M+R=25
J+(J+3)+(J+7)=25
3J+10=25
3J=15
J=5 ANSWER 1: Jack is 5 years old.
M=J+3=5+3=8 ANSWER 2: Martin is 8 years old.
R=J+7=5+7=12 ANSWER 3: Rob is 12 years old.
.
CHECK:
.
J+M+R=25
5+8+12=25
25=25